# V2X-R: Cooperative LiDAR-4D Radar Fusion with Denoising Diffusion for 3D Object Detection

### v3

@@ 7.3 Information of data collection
Split Seq Name Agent City
Train 1 2024_06_24_19_56_04 4CAVs Town06
2 2024_06_24_20_24_02 2CAVs Town06
3 2024_06_24_21_03_22 3CAVs Town06
4 2024_06_24_21_23_27 2CAVs Town06
5 2024_06_24_21_40_16 2CAVs Town06
6 2024_06_24_21_46_47 1CAVs; 1Infra. Town06
5 2024_06_25_15_28_10 2CAVs Town06
8 2024_06_25_15_46_09 4CAVs Town06
9 2024_06_25_16_28_47 2CAVs Town05
10 2024_06_25_17_02_14 2CAVs Town05
11 2024_06_25_17_12_23 2CAVs; 1Infra. Town05
12 2024_06_25_18_56_34 3CAVs; 1Infra. Town05
13 2024_06_26_09_51_48 4CAVs Town05
14 2024_06_26_10_31_36 2CAVs Town05
15 2024_06_26_10_44_26 3CAVs Town05
16 2024_06_26_11_05_36 3CAVs Town05
17 2024_06_26_19_17_09 3CAVs; 1Infra. Town05
18 2024_06_26_19_38_06 2CAVs; 1Infra. Town04
19 2024_06_26_20_18_14 2CAVs; 1Infra. Town05
20 2024_06_26_21_04_51 3CAVs; 1Infra. Town04
21 2024_06_26_21_15_34 2CAVs; 1Infra. Town05
22 2024_06_26_21_22_24 2CAVs; 1Infra. Town05
23 2024_06_26_21_29_14 1CAVs; 1Infra. Town04
24 2024_06_26_21_47_03 1CAVs; 1Infra. Town04
25 2024_06_26_21_56_21 2CAVs; 1Infra. Town10HD
26 2024_06_26_22_02_55 2CAVs; 1Infra. Town10HD
27 2024_06_26_22_12_38 4CAVs; 1Infra. Town10HD
28 2024_06_26_22_18_45 3CAVs Town10HD
29 2024_06_26_22_30_42 2CAVs Town07
30 2024_06_26_22_37_45 2CAVs Town07
31 2024_06_27_10_09_59 3CAVs Town07
32 2024_06_27_11_20_21 4CAVs Town03
33 2024_06_27_14_22_27 3CAVs; 1Infra. Town01
Validation 34 2024_06_28_17_45_09 3CAVs Town05
35 2024_06_28_18_48_16 2CAVs; 1Infra. Town05
36 2024_06_28_19_01_25 2CAVs; 1Infra. Town05
37 2024_06_28_19_20_09 2CAVs; 1Infra. Town05
38 2024_06_28_19_30_57 2CAVs; 1Infra. Town10HD
39 2024_06_28_19_45_02 5CAVs Town02
Test 40 2024_06_27_16_27_32 2CAVs Town06
41 2024_06_27_17_05_43 4CAVs; 1Infra. Town06
42 2024_06_27_17_41_33 3CAVs Town06
43 2024_06_27_19_09_42 5CAVs Town05
44 2024_06_28_09_17_57 2CAVs Town05
45 2024_06_28_09_33_41 3CAVs; 1Infra. Town05
46 2024_06_28_09_52_03 2CAVs; 1Infra. Town05
47 2024_06_28_10_07_17 2CAVs; 1Infra. Town04
48 2024_06_28_10_22_46 2CAVs; 1Infra. Town04
49 2024_06_28_10_58_01 3CAVs; 1Infra. Town10HD
50 2024_06_28_11_17_03 3CAVs; 1Infra. Town10HD
51 2024_06_28_14_59_30 4CAVs Town10HD
52 2024_06_28_15_26_15 3CAVs Town07
53 2024_06_28_15_52_12 2CAVs Town04
54 2024_06_28_16_18_18 2CAVs Town04
55 2024_06_28_16_18_18 4CAVs; 1Infra. Town03
Table 10: Information of data collection. CAVs denote Connected Automatic Vehicles and Infra. denotes Infrastructure.
@@
We summarize the critical collection details of each scene sequence by splits, as shown in Table 10, which includes agent information and scene sources for each sequence. Additional detailed acquisition information, such as the configuration of each agent’s trajectory, can be queried in the documentation of our V2X-R dataset.
8 Training Details of Benchmark
@@
AttFuse [54] L+4DR Sedan $AP_{BEV}$ 70.3 68.0 89.4 90.5 79.5 66.9 88.3 60.4
$AP_{3D}$ 69.0 66.8 79.4 88.6 70.7 59.2 86.2 58.6
Bus $AP_{BEV}$ 64.3 59.4 75.7 - 0.4 66.2 80.7 70.8
$AP_{3D}$ 51.0 53.3 75.6 - 0.2 65.6 76.8 36.3
$AP_{3D}$ 74.0 67.2 85.5 89.6 75.7 64.6 84.5 59.8
Bus $AP_{BEV}$ 64.1 55.3 72.0 - 15.1 62.7 97.5 73.9
$AP_{3D}$ 54.6 52.8 71.0 - 15.0 61.3 85.2 42.6
9 Configuration of Weather Simulation
@@ 10.1 K-Radar dataset and evaluation metrics
The K-Radar dataset [28] contains 58 sequences with 34944 frames of 64-line LiDAR, camera, and 4D radar data in various weather conditions. According to the official K-Radar split, we used 17458 frames for training and 17536 frames for testing. We adopt two evaluation metrics for 3D object detection: $AP_{3D}$ and $AP_{BEV}$ of the class ”Sedan” and ”Bus” at IoU = 0.3. We use the newest version (v2.1) of the label.
10.2 Performance on various weather.
11 Implementation of MDD
Component Parameter Value
Denoiser (U-net) input_channel 128
mid_channel 128
timestep_channel 64
output_channel 64
number_layers 2
number_resblock 2
Diffusion Process timesteps 3
betas [0.005,0.0275,0.05]
Table 14: The implementation details of our MDD module.
@@ 12 Visualization of V2X-R Dataset
Figure 6: Visualization of our V2X-R dataset and VoD [30] real-world dataset. The L+4DR in the last column indicates that the LiDAR point cloud is visualized together with the 4D radar point cloud, where to distinguish between them, we use colored dots (slightly smaller) for the LiDAR point cloud and orange dots (slightly larger) for the 4D radar point cloud. Color point clouds are assigned by z-axis values.

### v4

@@ 7.3 Information of data collection
We saved the critical collection details of each scene sequence by splits (can be found in "data_protocol.yaml" of the released dataset), which includes agent information and scene sources for each sequence. Additional detailed acquisition information, such as the configuration of each agent’s trajectory, can be queried in our V2X-R documentation.
8 Additional Results Analysis
Modality
Communication volumes (4B as unit, log-scale)/
Latency time (ms as unit, 27Mbps as transmission speed)
AttFuse V2XViT CoBEVT SICP L4DR
LiDAR-based 18.2/357.7 16.0/75.2 17.6/226.1 14.9/35.2 -
4D radar-based 14.6/28.4 14.0/18.6 15.0/37.4 13.6/14.7 -
LiDAR+4D radar 18.4/402.1 16.9/142.0 18.1/317.6 15.2/42.7 17.7/253.1
Table 10: Analysis of Communication Costs and Bandwidth for Different Models (SM2MM,SA2MA).
8.1 Transmission cost and bandwidth.
We have calculated transmission cost (count of non-zero elements in the feature map) and latency for different modalities. As shown in Table 10, 4D radar has the advantage of sparse feature transmission, and its fusion with LiDAR brings acceptable transmission cost and latency time.
Std of localization error (m)/(Val-AP@50)
Method 0.0 0.1 0.2 0.3 0.4 0.5 0.6
AttFuse 84.30 81.28 79.87 77.84 73.09 68.30 62.84
CoBEVT 87.02 85.26 83.30 81.33 77.83 70.79 63.47
AdaFusion 87.31 87.15 86.25 84.22 79.79 72.89 64.15
Where2comm 85.78 85.58 84.78 82.91 79.57 76.23 71.77
Table 11: Performance of different models under varying degrees of localization error.
AttFuse [57] L+4DR Sedan $AP_{BEV}$ 70.3 68.0 89.4 90.5 79.5 66.9 88.3 60.4
$AP_{3D}$ 69.0 66.8 79.4 88.6 70.7 59.2 86.2 58.6
Bus $AP_{BEV}$ 64.3 59.4 75.7 - 0.4 66.2 80.7 70.8
$AP_{3D}$ 51.0 53.3 75.6 - 0.2 65.6 76.8 36.3
$AP_{3D}$ 74.0 67.2 85.5 89.6 75.7 64.6 84.5 59.8
Bus $AP_{BEV}$ 64.1 55.3 72.0 - 15.1 62.7 97.5 73.9
$AP_{3D}$ 54.6 52.8 71.0 - 15.0 61.3 85.2 42.6
@@ 8 Additional Results Analysis
8.2 localization noise
Following Where2comm [12], we added different degrees of localization error to our V2X-R dataset to conduct experiments. As shown in Table 11, all methods experience performance degradation to varying degrees as localization error increases. This helps to explore performance under real localization errors.
8.3 Performance of various weather on K-Radar.
Component Parameter Value
Denoiser (U-net) input_channel 128
mid_channel 128
timestep_channel 64
output_channel 64
number_layers 2
number_resblock 2
Diffusion Process timesteps 3
betas [0.005,0.0275,0.05]
Table 14: The implementation details of our MDD module.
9 Training Detail
9.1 Benchmark
@@ 9.2 Implementation of MDD
9.3 Configuration of weather simulation
@@ 9.4 K-Radar dataset and evaluation metrics
The K-Radar dataset [31] contains 58 sequences with 34944 frames of 64-line LiDAR, camera, and 4D radar data in various weather conditions. According to the official K-Radar split, we used 17458 frames for training and 17536 frames for testing. We adopt two evaluation metrics for 3D object detection: $AP_{3D}$ and $AP_{BEV}$ of the class "Sedan" and "Bus" at IoU = 0.3. We use the newest version (v2.1) of the label.
Figure 6: Visualization of our V2X-R dataset and VoD [33] real-world dataset. The L+4DR in the last column indicates that the LiDAR point cloud is visualized together with the 4D radar point cloud, where to distinguish between them, we use colored dots (slightly smaller) for the LiDAR point cloud and orange dots (slightly larger) for the 4D radar point cloud. Colored point clouds are assigned by z-axis values.